# A Multi-Objective Evaluation Framework for Analyzing Utility-Fairness Trade-Offs in Machine Learning Systems

### v1

@@ IV Empirical Validation
System $HV$ $\widehat{ONVG}$ $ONVGR$ $\overline{UD}$ $OS^{*}$ $\widehat{\Delta}$
$System1$ 0.75 0.84 0.48 0.97 0.30 0.30
$System2$ 0.73 1.00 0.57 0.98 0.30 0.37
${}^{*}1e-2$
(c) Quantitative results for the systems as white-box.
Figure 13: Empirical study with the proposed evaluation framework comparing multidimensional utility-fairness trade-offs from two ML systems.
V Discussion
@@
This work presents the requirements and advantages of an evaluation framework assessing multidimensional utility-fairness trade-offs obtained with ML systems. The framework enhances the comparison of different modeling strategies, with the goal of selecting optimal solutions for real-world applications that require the assessment of multiple fairness criteria. The ease-of-use and effectiveness of the proposed framework are explored through comprehensive simulations and empirical studies using a medical image dataset designed to test fairness optimization approaches. Unlike previous evaluation frameworks, this work provides a series of steps for the selection process of ML systems in the context of multidimensional fairness exploring different criteria, supported by MOO principles. Characterizing the optimal PF is particularly useful in tasks where contradictory fairness performance indicators cannot be avoided, and trade-offs should be tuned to specific fairness requirements from decision-makers. Our framework could guide the tuning of ML systems to different notions of fairness for any sensitive attribute and metric in a simple and transparent manner.
@@
As mentioned in Section II, the comprehensive analysis of PFs in multiple dimensions is a limitation of previously proposed fairness evaluation approaches. A summarized objective representation of performance indicators from the PFs, both as a radar chart and as a measurement table, overcomes the limitations from visualizing performance plots in multiple dimensions resulting from the assessment of different ML systems. Furthermore, analogous to AUC for ROC, our evaluation framework can transform the qualitative trend into a quantitative measurement, thus gathering all necessary information for the interpretation of the performance gap between the trade-off systems in consideration, and against an ideal solution.
@@
There may be some limitations of the proposed evaluation framework. A restriction is the exponential cost for computing the MOO-based performance indicators as the number of objectives increases [37]. However, we argue that the majority of tasks studied in the literature focus on a small set of sensitive attributes, i.e., gender, age and race. The use of other performance indicators, such as Inverted Generational Distance [34], could be explored as alternative measurements to address this issue. Another issue to consider, is the equal weighting of all performance indicators, which normalizes the contribution of the different indicators in the final evaluation. This may not be desired in situations where an indicator can evaluate the systems in suboptimal performance and needs to have less impact on the final decision compared to others. The proposed evaluation framework could be extended to support the dynamic weighting of the indicators, so that the re-weighting can be performed based on the use-case. As a final limitation, there may also be situations in which the indicator measurements are the same or not feasible (refer OS and UD in Fig. 10, respectively) for all systems in comparison. In this case, such an indicator is useless to consider for the final decision and the proposed framework simplifies the evaluation to the joint contribution of the rest. This issue could be alleviated by using different indicators that discriminate more, as the proposed framework supports seamlessly including/excluding different types of indicators.
VI Conclusions
@@
This paper proposes a multi-objective evaluation framework for utility-fairness trade-offs resulting from ML systems, using performance indicators based on MOO. The proposed framework is model and task agnostic, allowing for high flexibility in the comparison of ML strategies, even when they have been optimized for different objectives. This is an adaptive assessment framework that supports any kind of performance indicators, including the proposed method, for convergence, diversity and capacity analysis. The proposed framework is able to perform a comprehensive analysis of ML systems with a measurement table and radar chart, overcoming the limitations resulting from the qualitative assessment of solutions with multiple fairness requirements. These tools provide a structured and visual means to evaluate and compare multiple fairness metrics in machine learning systems. The measurement table allows for a clear, organized presentation of the data, while the radar chart offers a visual representation of how well the system performs across various fairness criteria. The effectiveness of the evaluation approach is verified by performing simulations and empirical analysis for a variety of use-cases, both with black and white box ML models. The proposed system is made available for public access to be applied in the context of multi-objective evaluation.
VII Acknowledgement

### v2

@@ 4.3 Evaluation Results
System $\bm{\mathit{HV}}$ $\bm{\widehat{\mathit{ONVG}}}$ $\bm{\mathit{ONVGR}}$ $\bm{\mathit{UD}}$ $\bm{\mathit{AS}}$ $\bm{\widehat{\Delta}}$
System1 0.64 0.83 0.56 0.65 0.49 0.40
System2 0.70 1.00 0.43 1.00 0.39 0.44
(c) Quantitative results.
Figure 14: DenseNet/LoRA-ViT-Small on mBRSET.
(a) Pareto plot.
(b) Radar chart.
Convergence-Diversity Capacity Diversity
Distribution Spread
System $\bm{\mathit{HV}}$ $\bm{\widehat{\mathit{ONVG}}}$ $\bm{\mathit{ONVGR}}$ $\bm{\mathit{UD}}$ $\bm{\mathit{AS}}$ $\bm{\widehat{\Delta}}$
System1 0.86 1.00 0.15 0.68 0.22 0.29
System2 0.79 0.64 0.10 0.65 0.45 0.26
(c) Quantitative results.
Figure 15: DenseNet/LoRA-ViT-Small on Shenzhen.
The comparative evaluation of the DenseNet- and LoRA-ViT-Small-based systems on the mBRSET dataset is illustrated in Fig. 14, which reports the distribution of utility (F1-score) and fairness outcomes (equalized odds difference (EOD) for the obesity attribute) across all ML models in the trade-off system. We note that group-wise comparisons based on prevalence-sensitive metrics such as F1-score should be interpreted with caution, as they are influenced by the underlying base rate (class ratio) and may partially reflect distributional characteristics rather than purely model-induced disparities. The curves show that System2 has lower EOD scores for a broader range of utility levels, indicating improved fairness consistency under varying preference settings. Table 14c further summarizes the evaluation for PF. System2 achieves higher scores in both $\widehat{\mathit{ONVG}}$ and UD. The HV metric also favors System2 ($0.70$ vs. $0.64$), reflecting a closer alignment with the ideal reference point. In contrast, System1 exhibits a higher ONVGR and AS values. When aggregated using the area score $\widehat{\Delta}$, System2 achieves $0.44$ compared to $0.40$ for System1, confirming that System2 exhibits a higher overall trade-off performance structure when balancing diagnostic utility and obesity fairness on this dataset.
(b) System2 (orange) slightly dominates System1 (blue) with a margin.
Convergence-Diversity Capacity Diversity Model 1 Model 10 Model 25
Distribution Spread Utility Fairness Utility Fairness Utility Fairness
System $\bm{\mathit{HV}}$ $\bm{\widehat{\mathit{ONVG}}}$ $\bm{\mathit{ONVGR}}$ $\bm{\mathit{UD}}$ $\bm{\mathit{AS}}$ $\bm{\widehat{\Delta}}$ $\bm{\mathit{Acc}}$ $\bm{\mathit{F1}}$ $\bm{\mathit{DP}}$ $\bm{\mathit{EOdd}}$ $\bm{\mathit{Acc}}$ $\bm{\mathit{F1}}$ $\bm{\mathit{DP}}$ $\bm{\mathit{EOdd}}$ $\bm{\mathit{Acc}}$ $\bm{\mathit{F1}}$ $\bm{\mathit{DP}}$ $\bm{\mathit{EOdd}}$
System1 0.76 0.86 0.35 0.70 0.16 0.28 0.78 0.77 0.07 0.09 0.68 0.75 0.06 0.08 0.54 0.70 0.00 0.00
System2 0.73 1.00 0.46 0.77 0.15 0.35 0.76 0.76 0.08 0.10 0.56 0.71 0.02 0.05 0.54 0.70 0.00 0.00
(c) Quantitative results.
Figure 16: PHN on HGF.
The performance of the DenseNet- and LoRA-ViT-Small-based systems on Shenzhen dataset is summarized in Fig. 15 that visualizes the utility performance (F1-score) and fairness disparity (min-max difference for the gender groups). As shown in the plots, System1 demonstrates a more favorable balance between fairness variability and utility consistency. Specifically, System1 maintains a higher performance on the F1-score without hitting very low scores while achieving comparable EOD values against System2. A quantitative comparison using PF-based indicators is provided in Table 15c. System1 achieves higher performance in $\widehat{\mathit{ONVG}}$ with ONVGR, and outperforms System2 in UD and HV ($0.86$ vs. $0.79$). While System2 exhibits a larger AS indicating a broader spread of trade-off solutions, its distribution characteristic is weaker. Aggregating all indicators through the area score shows System1 achieving $0.29$ compared to $0.26$ for System2, validating that System1 provides a higher trade-off structure for gender fairness in the Shenzhen dataset.
For the HGF experiment, Fig. 16 summarizes the evaluation of the two PHN-based systems considering one utility performance (accuracy) and two fairness metrics (EOD for the gender and race). Visual inspection alone illustrated by the 3D performance plot in Fig. 16a is insufficient to reliably differentiate the systems when multiple fairness criteria are considered alongside utility. At this point, the radar visualization in Fig. 16b provides a clearer qualitative aspect by showing a slight performance improvement for System2 over System1. This observation is further confirmed by the quantitative indicators reported in Table 16c. Although there is an inconsistency between HV and the other performance indicators ($\widehat{\mathit{ONVG}}$, ONVGR and UD), an area of $0.35$ over $0.28$ shows that System2 constitutes the better-performing utility-fairness trade-offs than System1. In this case, it is not possible to differentiate the systems over AS as they both exhibit average spreads of $0.15$ and $0.16$. Finally, it is straightforward to interpret that both systems are far from optimality as they are well below the ideal performance of $\widehat{\Delta}=1.00$.
Table 16c also reports the performance of individual sub-NNs generated by the PHN (Models 1, 10, and 25), providing concrete examples of model-level behavior within each utility-fairness trade-off system. These examples illustrate how operating points guided by preference along the PF can differ substantially in their balance between utility and fairness. For instance, in System1 (gender fairness), the first sub-model achieves high utility (accuracy $\mathit{(Acc)}=0.78$ and F1-score $\mathit{(F1)}=0.77$), but exhibits demographic disparity (demographic parity difference $\mathit{(DP)}=0.07$ and equalized odds difference $\mathit{(EOdd)}=0.09$), whereas later models, such as 25th sub-NN, perform substantially reduced unfairness ($\mathit{DP}=0.00$ and $\mathit{EOdd}=0.00$) at the cost of lower accuracy and F1-score ($\mathit{Acc}=0.54$ and $\mathit{F1}=0.70$). A similar trend is observed for System2 (race fairness), where earlier preference vectors favor utility while later ones build more equitable outcomes across racial groups. However, evaluating these systems solely through individual model outputs is insufficient, as meaningful assessment requires understanding their behavior under multiple fairness criteria. Although System1 and System2 are trained and evaluated on gender and race fairness, respectively, a complete analysis also demands examining cross-criterion disparities to capture the broader fairness landscape. These examples highlight how the proposed framework supports aggregated, system-level interpretation rather than relying on isolated operating points. Ultimately, this reinforces that fairness-aware evaluation must adopt a multi-objective perspective that jointly reflects utility and multiple fairness criteria.
5 Discussion
@@
This work presents the requirements and advantages of an evaluation framework assessing multidimensional utility-fairness trade-offs obtained with ML systems. The framework enhances the comparison of different modeling strategies, with the goal of selecting optimal solutions for real-world applications that require the assessment of multiple fairness criteria. The ease of use and effectiveness of the proposed framework are explored through comprehensive simulations and empirical studies using medical imaging datasets designed to test fairness optimization approaches. Unlike previous evaluation frameworks, this work provides a series of steps for the selection process of ML systems in the context of multidimensional fairness exploring different criteria, supported by MOO principles. Characterizing the optimal PF is particularly useful in tasks where contradictory fairness performance indicators cannot be avoided, and trade-offs should be tuned to specific fairness requirements from decision-makers. Our framework could guide the tuning of ML systems to different notions of fairness for any sensitive attribute and metric in a simple and transparent manner. Fairness is interpreted as a multidimensional evaluation space, and the spectrum of trade-offs is captured across demographic attributes rather than enforcing a single optimal fairness configuration, which may fail to represent the overall capability of a fairness-aware ML system.
While the utility-fairness trade-off is commonly reported in fairness-aware ML studies, it does not universally manifest across all settings. As discussed in previous works (Chouldechova, 2017; Kleinberg et al., 2016), different fairness definitions such as independence, separation, and sufficiency are mutually incompatible under certain statistical assumptions, which means that optimizing multiple fairness notions simultaneously is impossible to achieve. Furthermore, recent works also show that the perceived utility-fairness tension may not always exist when dataset-level biases are addressed (Wick et al., 2019; Dutta et al., 2020). Conversely, a more recent work by Dehdashtian et al. (2024) suggests that utility-fairness trade-offs can indeed arise from intrinsic dataset characteristics by empirically observing the persistence of such tension. Taken together, these findings indicate that utility-fairness interactions are highly context dependent and influenced by both data distribution and model assumptions. In this ongoing research landscape, our model- and metric-agnostic evaluation framework contributes by providing a structured way to capture a snapshot of how different methods achieve the balance/compensation between utility and fairness. By quantifying these dynamics in a unified evaluation space, the framework allows for a systematic comparison of fairness-aware ML systems independent of their underlying architectures or optimization strategies, and it provides a baseline for future work that needs a benchmark for multi-objective fairness evaluation. Moreover, the framework is designed flexibly for practitioners so that they can select and evaluate the specific utility-fairness trade-offs relevant to their application, rather than imposing any pre-determined balance among objectives or assuming a universal trade-off structure.
@@
As mentioned in Section 2, the comprehensive analysis of PFs in multiple dimensions is a limitation of previously proposed fairness evaluation approaches. A summarized objective representation of performance indicators from the PFs, both as a radar chart and as a measurement table, overcomes the limitations from visualizing performance plots in multiple dimensions resulting from the assessment of different ML systems. Furthermore, analogous to AUC over ROC, our evaluation framework can transform the qualitative trend into a quantitative measurement, thus gathering all necessary information for the interpretation of the performance gap between the trade-off systems in consideration and against an ideal solution.
Beyond the medical imaging domain, the proposed framework has broader implications for fairness evaluation across ML applications, as our approach generalizes the evaluation process by integrating multiple fairness and utility criteria within a single multi-objective formulation, independent of the domain under consideration. This generalization enables the framework to serve as a common evaluation protocol for comparing utility-fairness trade-offs in different contexts, such as financial risk assessment and biometric systems. In this regard, our results complement existing fairness benchmarking protocols by providing a structured, quantitative, and qualitative means of summarizing utility-fairness trade-off performance across domains.
@@
There may be some limitations of the proposed evaluation framework. A restriction is the exponential cost for computing the MOO-based performance indicators as the number of objectives increases (Audet et al., 2021). This scalability challenge is particularly relevant when fairness must be evaluated across many demographic axes or multiple fairness notions simultaneously, which may arise in complex real-world deployments. However, we argue that the majority of tasks studied in the literature focus on a small set of sensitive attributes, i.e., gender, age, and race. The use of alternative performance indicators, such as Inverted Generational Distance (IGD) (Coello Coello and Reyes Sierra, 2004), could be explored to improve efficiency. Another issue to consider is the equal weighting of all performance indicators, which normalizes the contribution of the different indicators in the final evaluation. This may not be desirable in situations where an indicator can evaluate the systems in suboptimal performance and needs to have less impact on the final decision compared to others. The proposed evaluation framework could be extended to support the dynamic weighting of the indicators, so that the re-weighting can be performed based on the use case. As another limitation, there may also be situations in which the indicator measurements are the same or not feasible (refer to AS and UD in Fig. 10) for all systems in comparison. In this case, such an indicator is not informative for the final decision and the proposed framework simplifies the evaluation to the joint contribution of the rest. This issue could be alleviated by using different indicators that discriminate more, as the proposed framework supports seamlessly including/excluding different types of indicators. Finally, as emphasized by Selbst et al. (2019), algorithmic modeling and evaluation of fairness cannot be fully abstracted from its social context. Accordingly, the proposed evaluation framework should be interpreted as an assessment tool for examining utility-fairness trade-offs under clearly defined assumptions and constraints, rather than as a universal fairness solution applicable to all real-world scenarios.
6 Conclusions
@@
This paper proposes a multi-objective evaluation framework for utility-fairness trade-offs resulting from ML systems, using performance indicators based on MOO. The proposed framework is model- and task-agnostic, allowing for high flexibility in the comparison of ML strategies, even when they have been optimized for different objectives. This is an adaptive assessment framework that supports any kind of performance indicators, including the proposed method, for convergence, diversity and capacity analysis. The proposed framework is able to perform a comprehensive analysis of ML systems with a measurement table and radar chart, overcoming the limitations resulting from the qualitative assessment of solutions with multiple fairness requirements. These tools provide a structured and visual means to evaluate and compare multiple fairness metrics in ML systems. The measurement table allows for a clear, organized presentation of the data, while the radar chart offers a visual representation of how well the system performs across various utility and fairness criteria using MOO indicators. The effectiveness of the evaluation approach is verified by performing simulations and empirical analyses for a variety of use cases, with both black- and white-box ML systems. In particular, the empirical results on medical imaging based use cases illustrate how the framework can expose fairness disparities between diagnostic models, guiding the selection of appropriate trade-offs for ML systems in healthcare applications. The proposed system is made available for public access to be applied in the context of multi-objective evaluation for any domain.
[…]
@@
[…]
The work follows appropriate ethical standards in conducting research and writing the manuscript, following all applicable laws and regulations regarding treatment of animals or human subjects.
[…]
We declare we don’t have conflicts of interest.
[…]
All datasets used in this study are publicly available: